The fundamental notes produced in an open end organ pipe and closed end organ pipe of same length are in the ratio

Grade:11

1 Answers

Avinash Kumar
7 years ago
ans is 2.
for open organ pipe fundamental frequency is f1 = v/2l.
for closed organ pipe fundamental frequency is f2 = v/4l.
f1/f2=2